# Do you monitor CI Practices? I don’t know. You tell me: A case study

### v1

@@ Abstract.
Background: In this paper we seek for understand the benefits and challenges of monitoring CI practices in day-to-day software development. Aims: We aim to evaluate the impact of monitoring seven CI practices in a real-world scenario on three public organizations in Brazil. Method: We first developed a CI practices monitoring suite tool and conducted a multiple-case study applying a mixed-methods strategy, combining surveys, interviews, log data, and mining data from CI services. Results: We verified organization’ interest in monitoring CI practices. Monitoring provided an overview of the organization’s CI status, not covered by other tools, motivated constant improvement in these practices, a perception of software quality, improve the communication and and it is easy to adopt. Conclusions: We recommend that companies adopt monitoring of CI practices and that CI services integrate this monitoring into their dashboards.
CI Practices, Monitoring, Case Study, CI Maturity
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2024; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software maintenance toolsccs: Software and its engineering Software performanceccs: Software and its engineering Software reliability
1. Introduction
@@
Continuous Integration (CI) is a software development practice where team members integrate their work frequently (Fowler, 2006). The integration process involves an automated build step, which includes compiling the code and running automated tests. This ensures the detection of integration errors as quickly as possible. Despite of extensive research into the benefits and costs of Continuous Integration  (Laukkanen et al., 2015; Widder et al., 2019; Hilton et al., 2017; Cassee et al., 2020; Hilton et al., 2016; Ståhl and Bosch, 2013a; Elazhary et al., 2021; Vasilescu et al., 2015; Bernardo et al., 2018; Ghaleb et al., 2019; Bach et al., 2017). Most of studies focused in miner repositories and study the overall impact of CI. There is a lack of studies specifically addressing the monitoring CI practices. Do developers monitor CI practices properly, and would they like to monitor? What benefits do they receive from this monitoring? These are some of the questions we aim to answer in this paper.
@@
In their Systematic Literature Review, Eliezio et al. (Soares et al., 2021) identified a lack of criteria for defining the use of CI. They found that 42.5% of the primary studies did not apply or establish any criteria to determine whether a project uses CI. Among those that did use criteria, 56.25% of the studies relied on only one criterion. The most common criterion applied was the adoption of a CI service, such as Travis CI. In the same direction, other studies (Felidré et al., 2019) have already demonstrated that some CI projects do not follow all CI practices. That study found that: (i) 60% of the projects had infrequent commits; (ii) 85% of the projects had at least one broken build that took a long time to be fixed; and (iii) most projects had builds that executed for more than 10 minutes. A fact known as CI THEATER.
@@
In this context, this paper aims to understand the benefits of using and monitoring CI practices by assessing their utility in real-world software development projects. To achieve this goal, we conducted a multiple case study with mixed methods, combining survey, interviews and access log analysis. For each project, we applied an initial survey to understand the level of CI maturity and the team’s satisfaction with CI. This allowed us to better compare and understand the impact of monitoring CI practices on projects at the end of the study. Additionally, we developed a tool with an extensive suite of metrics that automatically calculates the values of seven CI practices, enabling developers to continually have access to these metrics and receive weekly alerts about CI practice evolution. The CI practices are describe bellow:
  • Commit Per Weekday/Commit Activity  (Felidré et al., 2019): Mean of the absolute number of commits according to the week day of the analyzed period.
  • Coverage  (Felidré et al., 2019): It measures which parts of a program are executed when running the tests. Represents the percentage of the program covered by tests.
  • Build Duration  (Felidré et al., 2019): It measures the duration of the build (build finished at timestamp - build started at timestamp).
  • Build Activity/Build-frequency  (Ståhl and Bosch, 2014): It is a unit interval (i.e., a closed interval [0,1]) representing the rate of builds across days, i.e, if builds were made every day in the period of CI, the value would be 1. If builds were made in half of the days, the value would be 0.5. If there were no builds, the value would be 0.
  • Build Health/Build Quality  (Soares et al., 2021): It is a unit interval representing the rate of build failures across days. If there were build failures every day, the value would be 0. if there were no build failures, the value would be 1.
  • Time to Fix a Broken Build  (Felidré et al., 2019): It consists of the median time in a period that builds remained broken. When a build breaks, we compute the time in seconds until the build returns to the “passed” status. If the CI period ends and the build did not return to the “passed” status, we consider the time since it was broken until the end of the period. When a period has no broken builds, the value would be 0.
  • Comments Per Change  (Thompson, 2017; Ståhl and Bosch, 2013b): Mean of the number of comments grouped by Merge Request or Issue. This metric seeks to measure the level of communication between the team.
@@
Over a period of 8 weeks, we conducted a series of interviews with team members to gain insights into the benefits and challenges of monitoring these CI practices. We applied this suite of metrics in three software development projects from different public Brazilian organizations. By collecting metrics from CI services, access log, surveys and interviewing organizational members, we answered the following research questions:
@@
  • RQ1: Are the developers satisfied with the project’s CI maturity levels? By surveying and interviewing members of the project, we assess whether developers are satisfied with the current level of CI in their project. This exploratory research question gives us insight into the need to monitor CI practices. Our findings indicate that $23\frac${2}{3}divide start_ARG 2 end_ARG start_ARG 3 end_ARG of the projects were satisfied with the CI maturity level before the case study, although they did not utilize any metric-driven assessment to evaluate CI maturity. Nevertheless, despite initial satisfaction, CI monitoring helped participants to notice several aspects of CI that needed to be improved.
@@
  • RQ2: What benefits does CI monitoring bring? By interviewing members of the project, we collected team members’ perceptions regarding the benefits of monitoring CI practices. Our findings show that participants identified several benefits of CI monitoring, including its ability to provide an overview of the CI stage, highlight problems related to CI, and serve as a reminder and motivator for CI improvement, among other advantages.
@@
  • RQ3: What are the challenges/disadvantages of monitoring CI practices on a daily basis? By interviewing members of the project, we collected team members’ perceptions regarding challenges faced in monitoring CI practices. There were no significant challenges directly related to monitoring CI practices and the use of the monitoring tool. External problems, such as deadline management, instability in the CI environment, and a lack of emphasis on the importance of CI, were the main challenges faced by projects.
@@
  • RQ4: Are developers interested in using the tool? By collecting the log of access to dashboard of our suite of metrics, without the knowledge of the participants, we assessed how frequently the tool was used during the case study by the development team, thus measuring real interest in monitoring CI practices. The record of access to the dashboard screen allows us to conclude that there was considerable access to the tool during the case study, even outside the interview days. Mainly by companies that already had a certain level of CI maturity.
@@
  • RQ5: What was the evolution of CI practices during the use of the tool? By comparing the evolution of CI practices throughout the study period, we gain insights into the role of monitoring in this evolution. We noticed a notable enhancement in CI practices with one project, but for the other two projects, due to the instability of the CI environment and the break for the end-of-year holiday, we cannot perceive a constant improvement.
2. Continuous Integration Practices Monitoring
@@
Wang et. al. (Wang et al., 2022) presented an empirical study to observe the impact of test automation maturity on product quality, test automation effort, and release cycles in the Continuous Integration (CI) context of open-source projects. They carried out a test automation maturity survey with the main contributors of 149 open-source Java projects. “Test automation maturity” was quantified as the total score from all the maturity-related questions in their survey. Their findings indicated that higher levels of test automation maturity were positively correlated with improved product quality (p-value=0.000624) and shorter release cycles (p-value=0.01891).
@@
Sallin et. al.  (Sallin et al., 2021) conducted a comprehensive literature review to assess the feasibility of automatically measuring the four DORA Metrics (GmbH, 2023). They discovered that no scientific literature had specifically investigated the automatic measurement of these metrics. Nevertheless, their examination of grey literature revealed 16 papers discussing various aspects of automatic measurement. Additionally, they conducted a case study involving a Scrum team comprising ten individuals to evaluate the benefits of automating the measurement of DORA metrics.
@@
Elazhary et. al. (Elazhary et al., 2021) investigated the extent to which three software organizations implement ten Continuous Integration (CI) practices defined by (Fowler, 2006). They explored the benefits these practices bring and the challenges encountered during their implementation. This inquiry was conducted through a multiple-case study with mixed methods, focusing on three small to medium-sized software-as-a-service organizations. The study revealed that four practices exhibited significant differences in adoption between organizations: (i) For “Maintain a single source repository” practice: dealing with merge conflicts had a higher priority than the unclear benefit of using a single source repository, (ii) For “Make the build self-testing” practice: the perceived work required for integration tests outweighs the perceived benefits, (iii) For “Keep the build fast” practice: doing comprehensive testing outweighs having shorter builds), (iv) For “Automate deployment” practice: enhancing security by enforcing deployment privileges for some organizations outweighs a sense of change ownership.
@@
This paper aligns with the previously cited works. While Wang et. al. (Wang et al., 2022) concentrated on a single CI practice, specifically test automation, this paper takes a broader, more exploratory approach by evaluating a larger group of practices. In contrast with the Sallin et. al.  (Sallin et al., 2021) work, this paper focus on scope of CI practices. Furthermore, the paper aimed to evaluate the advantages of monitoring these metrics, extending beyond the focus solely on automation. Perhaps the work most aligned with ours is the work proposed by Elazhary et. al.(Elazhary et al., 2021). However, our analysis focuses on different CI practices. Additionally, a dashboard tool was developed as part of this paper, which automates the collection and visualization of the CI practices studied, thereby facilitating the adoption of monitoring CI practices. In our study, we implemented a metric suite tool that automatically collects and delivers 7 CI practice values to developers. This represents a significant advantage, particularly for the organization undergoing the case study. Our primary focus lies in assessing the benefits and challenges associated with the monitoring of CI practices over time. In contrast, Elazhary et. al. (Elazhary et al., 2021) emphasis is on exploring the benefits and challenges of implementing CI practices.
3. Methodology
@@
We conducted a multiple-case study (Stake, 2013) on monitoring CI practices with three software projects from different organizations. The main inclusion criteria for organizations in this study were: (i) location – we opportunistically chose organizations where we could interact with the developers and personally visit the organization to give more confidence in analyzing their data; (ii) availability – organizations had to provide a basic infrastructure to run our tool and allocate some of their developers time for interviews and validation of our findings; and (iii) the organization should have at least one project that should be using a CI service and generating regular builds.
The characteristics of selected organization are:
@@
  • Department of IT at a Public University (Organization A): Organization A is a public educational institution, specifically a department of Information Technology (IT) within a prominent public university in Brazil. It encompasses a wide range of academic and administrative IT projects, such as: a job vacancy system, “technology park” management system, access control system, website, etc. The university is known for its commitment to research and technological advancements, making it an ideal candidate for studying CI practices.
  • Software Project at a University with a Global Technology Company (Organization B): Organization B represents a unique collaboration Organization Between a Brazilian university and a global technology company. This software project focuses on cutting-edge technology and innovation in 5G networks. The involvement with a global technology company brings a distinct perspective on CI practices and the integration of industry and academia.
@@
  • State House of Representatives (Organization C): Organization C represents a branch of the state government responsible for legislative activities. As a public institution, the State House of Representatives plays a vital role in shaping legislation and policies at the state level in Brazil. The development team work on developing the project’s management software, such as: human resources department system, website, among others. The inclusion of this organization allows for a diverse perspective on CI practices, different from the academic environment.
@@
Because of the ease of obtaining access authorizations, Organization A was the first to initiate the interviews of case study, starting the interviews in November 2023 and concluding them in December 2023. We took advantage of this to use this organization as a pilot for our case study, as suggested by best practices (Yin, 2009). Organization B and C started the case study interviews in mid-December 2023 and finished it in early February 2024. We collected CI practice data from October 1, 2023, until the end of the interview period. This resulted in the evolution of practices being tracked over a period of 3 months for Organization A and 4 and a half months for Organizations B and C.
@@
Coincidentally, all three organizations participating in the case study utilized the same CI service, Gitlab. They employed Gitlab On-Premise, installed on the organizations’ Intranet, requiring specific access requests for each entity. To gather CI practices, we mined the build history of projects in the Gitlab CI service. Additionally, we collected the coverage metric from the SonarQube tool whenever the project used it (Organizations A and B utilized this tool, while Organization C did not generate these metric). We selected key members of the development team as study participants, considering developers with experience and basic knowledge of CI and the company’s development environment. Figure 1 provides an overview of all the steps involved in our data collection plan.
Figure 1. Data collection plan
@@
The data collection plan for this case study involves eight-step approach. We start with an initial presentation to introduce the case study’s objectives, importance, needs and potential benefits for projects that participate in it. This step was important to convince organizations to participate in the case study. Following the initial presentation, the project’s representative are asked to sign privacy declarations to ensure data confidentiality and obtaining permission from the projects to proceed with the study. We apply an initial survey (explained in detail in Section 3.1) to measure the project current CI maturity level perception for all study participants. Subsequently, training sessions are conducted to familiarize the teams with the developed metric suite tool for CI monitoring. The tool is then installed in the Intranet of organization, on a server provided by them. This makes it easier for participants to access our tool, which was available 24 hours a day in their work environment. The suite metric tool is then configured to start the weekly collection of CI practices. Finally, we perform weekly interviews with the development teams (explained in detail in Section 3.2) to capture real-time insights to measure the impact of CI practices monitoring.
@@
In conducting the case study, we adhered to a systematic and structured methodology based on (Tellis et al., 1997). 1) The first phase involved the design of a comprehensive case study protocol. This protocol, should determinate the required skills for case study and should be developed and reviewed. 2) The second phase centered on the actual execution of the case study. Adequate preparation for data collection, installation of CI practices monitoring suite and the distribution of questionnaires to relevant participants within the selected projects to assess the perception of current CI maturity in the project. Subsequently, interviews were conducted weekly to gather qualitative insights focused on answering the study’s research questions. 3) The third phase focused on the meticulous analysis of the gathered evidence. 4) The fourth phase involved the synthesis of conclusions, recommendations, and implications based on the evidence obtained through the case study. We also following the best practices defined by (Yin, 2009): 1) Conducting a pilot case study (Organization A), 2) training for a specific case study, 3) screening candidate cases, 4) Ask good questions, 5) Be a good “listener”, 6) Stay adaptive, 7) Have a firm grasp of the issues being studied and 8) Avoid biases by being sensitive to contrary evidence.
@@
To encourage participation in the study, we also outline the benefits for projects participating in the study, they are: (i) the project will be able to have a overview vision of how its CI maturity process, (ii) The monitoring suite tool, along with its source code, was made available permanently and free of charge to the project, and (iii), at the end of the study, 10 more metrics will be unlocked, including Continuous Delivery and DevOps metrics (DORA metrics  [^1] https://cloud.google.com/blog/products/devops-sre/using-the-four-keys-to-measure-your-devops-performance) .
3.1. Survey

### v2

# Monitoring Continuous Integration Practices in Industry: A Case Study

@@ Abstract
In this paper, we study the benefits and challenges of monitoring Continuous Integration (CI) practices in software development. Our aim is to evaluate the impact of monitoring seven CI practices in industry using three organizations in Brazil as case studies. We developed a tool for monitoring CI practices and conducted a multiple case study, applying a mixed-methods strategy. We combined surveys, interviews, log data, and repositories data from software projects and their CI services.We gauged the organization’s interest in monitoring CI practices. The act of monitoring CI provided an overview of the organizational state of practice in terms of CI, motivated further improvement of CI practices, increased perceived quality of software, and improved project communication. We recommend that companies adopt the practice monitoring of CI practices and that CI services integrate monitoring functionalities into their dashboards.
Index Terms:
CI Practices, CI Metrics, Monitoring, Case Study, CI Maturity
I Introduction
@@
Continuous Integration (CI) is a software development practice in which team members frequently integrate their work [1]. The integration process involves an automated build step that includes compiling the code and running automated tests. This ensures the detection of integration errors as quickly as possible. Despite extensive research on the benefits and costs of Continuous Integration [2, 14, 11, 8, 12, 3, 9, 13, 7, 10, 4, 5], most studies focused on mining repositories to investigate the overall impact of CI. Although understanding the impact of CI is important, as a software engineering community, we do not know (1) whether development teams monitor their CI practices; (2) whether there exists willingness to monitor CI practices; and (3) which benefits the act of monitoring practices potentially bring to the development process. The focus of our research is to address these gaps.
@@
In their systematic review of the literature (SLR), Eliezio et al. [15] identified a lack of criteria to determine whether projects under study use CI properly. They found that 42.5% of the primary studies did not apply or establish any criteria to determine whether a project uses CI properly. Among those which established some criteria, 56.25% of them relied on a single criterion. The most common criterion applied was the adoption of a CI service, such as Travis CI. In a similar vein, other studies [16] have demonstrated that several projects using CI services do not employ all CI practices [5]. For example, 60% of the projects practiced infrequent commits; (ii) 85% of the projects contained at least one broken build which took a long time to be fixed; and (iii) most of the projects contained builds which lasted more than 10 minutes to be generated.
@@
Considering that projects may have difficulties to implement the best practices of CI, our paper investigates whether the act of monitoring CI practices can foster a better CI environment in industry. To achieve this goal, we conducted a multiple case study with mixed methods, combining survey, interviews and access log analysis. For each project, we applied an initial survey to understand the level of CI maturity and the team’s satisfaction with CI. This allowed us to better compare and understand the impact of monitoring CI practices on projects at the end of the study. Additionally, we developed a tool [6, 5] with an extensive suite of metrics that automatically calculates the values of seven CI metrics (related to CI practices we aim to monitor), enabling developers to continuously access these metrics and receive weekly alerts about their evolution.
@@
Over a period of 8 weeks, we conducted a series of interviews with team members to gain insights into the benefits and challenges of monitoring the investigated CI practices. We applied the suite of metrics in three software development projects from different Brazilian organizations. By collecting metrics from CI services, access log, surveys and interviewing organizational staff, we answered the following research questions:
@@
  • RQ1: Are developers satisfied with the project’s CI maturity levels? By surveying and interviewing members of the project, we assess whether developers are satisfied with the current level of CI in their project. This exploratory research question provides insights into the need to monitor CI practices. Our findings indicate that $23\frac${2}{3}divide start_ARG 2 end_ARG start_ARG 3 end_ARG of the projects were satisfied with the CI maturity level even before we conducted our case study. However, developers did not use any metric-driven assessment to evaluate the CI maturity at this stage. Despite initial satisfaction levels, CI monitoring helped participants notice several aspects of CI that required improvement.
@@
  • RQ2: What benefits does CI monitoring bring? By interviewing project staff, we collected their perceptions regarding the benefits of monitoring CI practices. Our findings reveal that participants identified several potential benefits from monitoring CI, including the ability to obtain an overview of the current practice with respect to CI, to highlight problems related to CI, and serve as a reminder and motivator for improving and practicing CI, among other advantages.
@@
  • RQ3: What are the challenges/disadvantages of monitoring CI practices on a daily basis? By interviewing project staff, we collected their perceptions of potential challenges and disadvantages related to monitoring CI practices. There were no significant challenges directly related to monitoring CI practices nor the use of our monitoring tool. External problems, such as deadline management, instability in the CI environment, and a lack of emphasis on the importance of CI, were the main challenges faced by projects.
@@
  • RQ4: Are developers interested in using our metrics suite tool? By collecting the logs of access to the dashboard of our metrics suite tool (which contains the suite of metrics), and without the knowledge of the participants, we assessed how frequently the tool was used during the case study, thus measuring real interest in monitoring CI practices. The record of access to the dashboard screen allows us to conclude that there was considerable access to the tool during the case study, even outside the interview days. This was mainly the case for companies that already hold a certain level of CI maturity.
@@
  • RQ5: How did CI practices evolve during the use of our metrics suite tool? By comparing the evolution of CI practices throughout the study period, we gain insights into the role of monitoring in this evolution. We noticed a notable enhancement in CI practices with one project, but for the other two projects, due to the instability of the CI environment and the break for the end-of-year holiday, we cannot perceive a constant improvement.
II Related Work
In this section, we present an overview of research works related to monitoring of CI practices.
@@ II Related Work
Wang et. al. [17] conducted an empirical study to observe the impact of test automation maturity on product quality, test automation effort, and release cycles in the context of CI in open-source projects. They conducted a test automation maturity survey with developers from 149 open-source Java projects. “Test automation maturity” was quantified as the total score obtained in all maturity-related questions in their survey. Their findings indicated that higher levels of test automation maturity were positively correlated with improved product quality (p-value=0.000624) and shorter release cycles (p-value=0.01891).
@@
Sallin et. al.  [18] conducted a comprehensive literature review to assess the feasibility of automatically measuring the four DORA Metrics [19]. They found that the scientific literature has not investigated the automatic measurement of DORA metrics. Their examination of grey literature revealed 16 articles discussing various aspects of automatic measurement. The authors also conducted a case study involving a Scrum team comprising ten individuals to evaluate the benefits of automating the measurement of DORA metrics.
@@
Elazhary et. al. [9] investigated the extent to which three software organizations implement ten Continuous Integration (CI) practices defined by [1]. They explored the benefits these practices bring and the challenges encountered during their implementation. This inquiry was conducted through a multiple case study with mixed methods, focusing on three small to medium-sized software-as-a-service organizations. The study revealed that four practices exhibited significant differences in adoption between organizations.
@@
Differently from previous work, our research focuses on the monitoring of CI, which consists of broader practices, such as short build durations and test coverage. For example, although Wang et. al. [17] focused on a single CI practice, specifically test automation, we evaluate the monitoring of a larger group of practices. In contrast with the Sallin et. al.’s  [18] work, we focus on evaluating the advantages of monitoring CI practices, extending beyond the scope of sole automation. Perhaps the work most aligned with ours is the work proposed by Elazhary et. al.[9]. However, our analysis focuses on different CI practices. Moreover, different from all previous work, we developed a metrics suite tool [6, 5] that enabled us to perform case studies, collecting and visualizing the CI practices of three software development companies. The implemented tool suite automatically collects and delivers 7 CI metrics values to developers. This represents a significant advantage, particularly for the organization undergoing the case study. Our primary focus lies in assessing the benefits and challenges associated with monitoring CI practices over time. In contrast, Elazhary et. al. [9] emphasis was on exploring the benefits and challenges of implementing CI practices.
III Methodology
@@
We conducted a multiple case study [20] on monitoring the CI practices of three software projects from three different organizations. The main inclusion criteria for organizations in this study were: (i) location – we opportunistically selected organizations where we could interact with developers and personally visit the organization yielding more confidence in our data analysis; (ii) availability – organizations had to provide a basic infrastructure to run our tool and allocate some of their developer time for interviews and validation of our findings; and (iii) the organization had to have at least one project that used a CI service and generating regular builds.
The characteristics of the selected organization are:
@@
  • Department of IT at a Public University (Organization A): Organization A is a public educational institution, specifically a department of Information Technology (IT) within a prominent public university in Brazil. It encompasses a wide range of academic and administrative IT projects, such as: a job vacancy system, a technology park management system, an access control system, a website, etc. The university is known for its commitment to research and technological advancements, making it an ideal candidate for studying CI practices.
  • Global Technology Company (Organization B): Organization B is a global technology company with a local development team based in our city, which allowed its inclusion in our case study. The software project monitored in our case study focuses on cutting-edge technology and innovation in 5G networks, offering a distinct perspective on CI practices from an industry standpoint.
@@
  • State House of Representatives (Organization C): Organization C is a branch of the government of a Brazilian state, which is responsible for legislative activities. As a public institution, the State House of Representatives plays a vital role in shaping legislation and policies at the state level in Brazil. The development team works on developing the project’s management software, such as: the human resources department system, website, among others.
@@
Due to ease of access, Organization A was the first to initiate interviews of our case study, starting the interviews in November 2023 and concluding them in December 2023. We used this organization as a pilot for the case study, as recommended by best practices [21]. Organizations B and C began their interviews in mid-December 2023 and completed them in early February 2024. We collected data related to CI practices from October 1, 2023, until the end of the interview period. This allowed us to track the evolution of practices over a period of three months for Organization A and four and a half months for Organizations B and C.
@@
Coincidentally, all three organizations participating in the case study utilized the same CI service: Gitlab. They employed Gitlab On-Premise, installed on the organizations’ intranets, which required specific access requests for each entity. To gather metrics related to CI practices, we mined the build history of projects in the Gitlab CI service. Additionally, we collected code coverage metrics from the Sonarqube tool whenever available (Organizations C did not use Sonarqube). We selected key members of the development team as study participants. The criteria for selecting these members included development experience, basic knowledge of CI and familiarity with the company’s development environment. Figure 1 provides an overview of all the steps involved in our data collection plan.
Figure 1: Data collection plan
@@
The data collection plan for the case study consists of an eight-step approach. It begins with an initial presentation to introduce the case study’s objectives, importance, needs, and potential benefits for the participating projects. This step was important to convince the organizations to participate in the case study. Following the initial presentation, project representatives were asked to sign privacy declarations to ensure data confidentiality and to obtain permission to proceed with the study. We applied an initial survey (explained in detail in Section III-A) to assess the current perceived maturity level of CI practices in each project. Subsequently, training sessions were conducted to familiarize the teams with the developed metric suite tool for CI monitoring. The tool was then installed on a server within the organization’s intranet, provided by the organization. This setup made it easier for participants to access the tool, which was available 24 hours a day in their work environment. The metric suite tool was then configured to start the weekly collection of CI practice data. Finally, we conducted weekly interviews with the development teams (explained in detail in Section III-B) to capture real-time insights and measure the impact of monitoring CI practices. The CI metrics collected during our case study are described below:
  • Commit Per Weekday/Commit Activity  [16]: Mean of the absolute number of commits according to the week day of the analyzed period.
  • Coverage  [16]: Measures which parts of a program are executed when running the tests. Represents the percentage of the program exercised by tests.
  • Build Duration  [16]: Measures the duration of the build (build finished at timestamp - build started at timestamp).
  • Build Activity/Build-frequency  [27]: This metric is a unit interval (i.e., a closed interval [0,1]) that represents the rate of builds generated during the CI period.
  • Build Health/Build Quality  [15]: This metric is a unit interval that represents the rate of build failures during the CI period.
  • Time to Fix a Broken Build  [16]: Measures the length of time (the mean time) that builds were left broken. When a build breaks, we computed the time in seconds until the build returned to the “passed” status.
  • Comments Per Change  [22, 3]: Measures the mean number of comments in a group of Merge Requests or Issues. This metric measures the level of communication between the team.
@@
In conducting the case study, we adhered to a systematic and structured methodology based on [23]. 1) The first phase involved the design of a comprehensive case study protocol. This protocol should determine the required skills for the case study and should be developed and reviewed. 2) The second phase centered on the actual execution of the case study, including adequate preparation for data collection, installation of the CI practices monitoring suite, and distribution of questionnaires to relevant participants within the selected projects to assess the perception of current CI maturity in the project. Subsequently, interviews were conducted weekly to gather qualitative insights focused on answering the study’s research questions. 3) The third phase focused on the meticulous analysis of the gathered evidence. 4) The fourth phase involved the synthesis of conclusions, recommendations, and implications based on the evidence obtained through the case study. We also adopted the best practices defined by [21]: conduct a pilot case study (Organization A), train for a specific case study, screen candidate cases, ask good questions, be a good “listener,” stay adaptive, have a firm grasp of the issues being studied, and avoid biases by being sensitive to contrary evidence.
@@
To encourage participation in the study, we also outlined the benefits for projects participating in the study. These benefits are: (i) the project would gain an overview of its CI maturity and processes, (ii) the monitoring suite tool, along with its source code, would be made permanently available and free of charge to the project, and (iii) at the end of the study, 10 additional metrics would be unlocked, including Continuous Delivery and DevOps metrics (DORA metrics [^1] https://cloud.google.com/blog/products/devops-sre/using-the-four-keys-to-measure-your-devops-performance).
III-A Survey